OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

I got this much so far, is it right? u=sqrt. 2^2+-4^2 v=sqrt.3^2+-8^2 u=4.5 v=8.5 dot=38.25

OpenStudy (irishboy123):

look at your dot again

OpenStudy (anonymous):

isnt dot multiplying u and v?

OpenStudy (irishboy123):

2x3 + (-4)x(-8)

OpenStudy (irishboy123):

=?

OpenStudy (anonymous):

so 38?

OpenStudy (irishboy123):

yes

OpenStudy (anonymous):

okay and now what do i do?

OpenStudy (michele_laino):

better is if we write the lengths of our vectors, like this: \[\Large \begin{gathered} \left\| u \right\| = \sqrt {4 + 16} = \sqrt {20} \hfill \\ \left\| v \right\| = \sqrt {9 + 64} = \sqrt {73} \hfill \\ \end{gathered} \]

OpenStudy (irishboy123):

\(\large cos \theta = \frac{\vec u \bullet \vec v}{|\vec u||\vec v|} \)

OpenStudy (irishboy123):

find arccos use calculator!

OpenStudy (anonymous):

Im so confused lol what do I put for cos theta= like what goes in the u and v?

OpenStudy (irishboy123):

calculate the RHS first then, ie using your numbers: \(\large \frac{38}{4.5 \times 8.5}\) what do you get we NEED TO optimise later but do this for now

OpenStudy (anonymous):

.9934640523 arccos=6.55

OpenStudy (irishboy123):

yeah! that's a correct answer for "your" numbers but as @Michele_Laino pointed out, your rounding is going to let you down use these instead \(\huge \frac{38}{\sqrt{20}\times \sqrt{73}}\) your calculator will now compute these square roots for you to a very high dgeree of accuracy, and you will get a much more accurate answer

OpenStudy (anonymous):

I got .9945054529

OpenStudy (irishboy123):

yep that should give a very accurate answer if you do not round yourself into oblivion! what do you get for \(\theta\) now?

OpenStudy (anonymous):

What how do I find theta again?

OpenStudy (irishboy123):

you say \(cos \theta = .9945054529\) so ask your calculator for the inverse cosine of \( .9945054529\) like you did last time....

OpenStudy (anonymous):

.1023736009

OpenStudy (irishboy123):

the issue here might be calculator this is what mine looks like and it is a fairly basic one

OpenStudy (irishboy123):

see how all information is in the calculator, so i am not rounding anything. that guarantees accuracy

OpenStudy (irishboy123):

i've also inadvertently given you the answer (!!) .... but anyways compare that to the 6.55 you originally calculated. the question wants you to be accurate to within 1/10 th of a degree, so 6.55 would have failed...

OpenStudy (anonymous):

Yeah whoa accuracy haha yeah my calculator doesn't read what yours reads I wonder why

OpenStudy (anonymous):

thank you so much can you help me in a last question though?

OpenStudy (irishboy123):

sure but make a new thread for each quetsion

OpenStudy (anonymous):

okie dokie!